a baloon moves vertically such stone is projected with horiz velocity u relative 2 baloon the stone always hits a river at a fixed distance 2u^2/g horiz away frm it find height of balloon as func of time

Grade:12

1 Answers

Ramesh V
14 years ago

assume ballon hz. velocity be 0

so stone hz vel. is u

let t be time taken to hit river

s = u*t so, t = s/u = 2u2/gu

height of ballon = h = 0*t - 1/2*g*t2

                             = -1/2*g*(4u4/u2*g2)

                           h  =  - 2u2/g

--

regards

Ramesh